# Toward a Team of AI-made Scientists for Scientific Discovery from Gene Expression Data

### v2

@@ 5.2.3 Performance of Data Preprocessing
Performance Table 3 illustrates the Data Engineer’s ability on preprocessing sub-task. Among different iterations, we achieve the optimal outcomes of 95.20% Success Rate, 34.60% Precision, and 36.00% Recall on single-step regression problems and 83.50% Success Rate, 11.27% Precision and 11.37% Recall on two-step regression task.
Different maximum number of reviews We also evaluate the performance of different review iterations on the data preprocessing task. Table 3 reveals that additional review iterations of the preprocessing task generally improve the overall Success Rate, Precision, Recall, and Jaccard index. Notably, within the single-step task, a second review significantly boosts the Success Rate to 95.20%, Precision to 34.60%, Recall to 36.00%, and the Jaccard index to 23.90%.
@@ Appendix A Details of the roles
Statistician Statistician agents are assigned the task of performing regression analysis on the preprocessed datasets delivered by the Data Engineer. Its objective is to identify the genes that are predictive of the disease status considering different conditions. To this end, two Statistician agents are designed to perform two categories of regression, i.e. single-step and two-step regression, as detailed in Section B. Both types of Statistician agents will follow the instruction derived from Section B to perform the regression, then analyze the output and interpret and report the results.
@@
Appendix B Details of Regression Analysis on Gene Expression Data
Notations We consider $\mathbf{X}\in\mathbb{R}^{n\times p}$ as the data matrix, where $n$ represents the number of samples, and $p$ the number of gene expressions. We will use $\mathbf{y}\in\mathbb{R}^{n}$ to denote the disease status. The basic data generation process is assumed to be
$\mathbf{y}=\mathbf{X}\beta+\epsilon,$ (1)
where we use $\beta$ to denote the coefficients to estimate and $\epsilon\sim N(\mathbf{0},\mathbf{I}\sigma_{\epsilon}^{2})$.
Vanilla Case In the most vanilla case, Lasso is chosen for its effectiveness in variable selection within high-dimensional datasets, isolating the most informative genes while controlling for overfitting. Lasso Tibshirani (1996) addresses this task by solving:
$\min_{\beta}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\beta\|_{2}^{2}+\lambda\|\beta\|,$ (2)
where $\beta$ is the coefficients to estimate (non-zero coefficients mean that Lasso selects these genes), and $\lambda$ is a hyperparameter.
Confounding Factor Detection The above Vanilla case is applied when there are no confounding factors in the data. However, before that, we need first check the data to inspect whether there are confounding factors. Inspired by the Marchenko-Pastur Law Johnstone (2001) and following a previous heuristic Wang et al. (2017), we inspect the “gap” of eigenvalues of the covariance matrix $\mathbf{X}\mathbf{X}^{T}$. If there is a clear gap in the eigenvalues, we consider the data are from multiple distributions, and thus, potentially with confounding factors. More formally, the confounding factor detection will return true if $e_{i}-e_{i+1}>1/n$ holds for any $i=0,1,\dots,t$, where $e_{i}$ stands for the $i$th eigenvalues if eigenvalues have been sorted descendingly, and $t$ stands for the maximum number of eigenvalues we inspect.
After the detection of confounding factors, if the agent believes there are no confounding factors in the data, it will directly follow the vanilla case above. However, if there are confounding factors, the agent will need to decide one of the following two approaches.
  • Regression with Confounding Factor Correction
  • Regression after Confounding Factor Correction
Regression with Confounding Factor Correction The first choice is to take advantage of the impressive empirical power of linear mixed model in correct confounding factors for genetic association studies Yu et al. (2006); Lippert et al. (2011); Wang et al. (2022b). With the setup of linear mixed model, we can formulate the problem with
$\mathbf{y}=\mathbf{X}\beta+\mathbf{Z}\mathbf{u}+\epsilon,$ (3)
where $\mathbf{Z}$ is an indication of the underlying population, and $\mathbf{u}\sim N(\mathbf{0},\mathbf{I}\sigma_{u}^{2})$ and $\epsilon\sim N(\mathbf{0},\mathbf{I}\sigma_{\epsilon}^{2})$, the dimension of $\mathbf{0}$ and $\mathbf{I}$ can be inferred from the context, and $\sigma$ denotes the standard deviation of the Gaussian distribution.
We use several tricks to estimate $\beta$ and the solution for a single column of $\mathbf{X}$ (denoted by $\mathbf{x}$) is
$\beta_{\mathbf{x}}=\dfrac{\mathbf{x}^{T}(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-1}\mathbf{y}}{\mathbf{x}^{T}(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-1}\mathbf{x}},$
because the trick will use $\mathbf{X}$ to replace the unknown $\mathbf{Z}$ Lippert et al. (2011); Wang et al. (2022b). Also, we use $\widehat{\delta}$ to denote the estimated ratio of the two $\sigma$.
Note that linear mixed model will estimate the coefficients individually for each gene, which is not necessarily ideal for our analysis. As an alternative choice, Statistician can also use the following.
Regression after Confounding Factor Correction In order to have a multivariate method, where the coefficients of genes can be considered together, the Statistician can also choose to correct the confounding factors first. The correction is mostly built upon the theory of how linear mixed model helps to correct confounding factors Wang et al. (2022b), and achieved by the rotation of $\mathbf{X}$ as following:
$\tilde{\mathbf{X}}=(\widehat{\delta}\mathbf{I}+\mathbf{X}\mathbf{X}^{T})^{-\frac{1}{2}}\mathbf{X}$
where $\widehat{\delta}$ is again the estimation ratio of the two sigmas.
With this rotation, one can perform Lasso in the vanilla case again with $\tilde{\mathbf{X}}$ and $\mathbf{y}$ for estimation of coefficients.
Incorporation of Conditions To make sure the regression analysis will incorporate different conditions such as age and gender, we will update our basic assumption of the data generation of Eq. 1 into
$\mathbf{y}=\mathbf{X}\beta+\mathbf{C}\alpha+\epsilon$ (4)
where we use $\mathbf{C}$ to denote additional conditions such as age and gender, and use $\alpha$ to denote the corresponding coefficient effect sizes.
Note that Eq. 4 is to replace Eq. 1 as an underlying data generation assumption. All the above solutions to estimate $\beta$ for Eq. 1 still apply to Eq. 4 by incorporating $\mathbf{C}$ into Eq. 2 or Eq. 3.
Two-step Regression While the above process offers a solution to incorporate the conditions when they are known. There are many cases that we do not necessarily have the conditions in the data. In this case, we will follow the following procedure, which we name two-step regression. It uses common genes between datasets that are known for their association with the condition, to estimate the conditions that our analyses need but do not know.
For example, we are interested in studying the association between disease $d$ and gene expressions given conditions $c$. However, there are no gene expression samples collected with both $d$ and $c$ appear in the same dataset. In this case, we will identify two datasets: Dataset 1 has $\mathbf{y}^{(1)}$ of disease $d$ and the corresponding gene expression $\mathbf{X}^{(1)}$, and Dataset 2 has $\mathbf{y}^{(2)}$ of the condition $c$ and the corresponding gene expression $\mathbf{X}^{(2)}$. We also need to identify a common set of genes $\mathbf{X}_{c}^{*}$ that appear in both datasets. In addition, the selection criteria for $\mathbf{X}_{c}^{*}$ also include their established biomedical relevance to the condition.
In our two-step regression, the first step is to construct the following regression model
$\mathbf{y}^{(2)}=\mathbf{X}^{(2)}_{c}\beta_{c}+\epsilon_{c}$ (5)
$\beta_{c}$ are the coefficients indicating the effects of these genes on the condition, and $\epsilon_{c}$ is the error term.
In the second step, we first predict the missing conditions for the first dataset, with
$\mathbf{C}=\mathbf{X}^{(2)}_{c}\widehat{\beta}_{c},$
where $\widehat{\beta}_{c}$ is the estimated $\beta_{c}$ from the first step.
Then, we create a $\mathbf{X}^{(2)}_{t}$ by excluding the column of $\mathbf{X}^{(2)}_{c}$ out of $\mathbf{X}^{(2)}$. Then the remaining follows by performing the analysis with $\mathbf{y}$, $\mathbf{X}^{(2)}_{t}$, and $\mathbf{C}$.
Appendix C Criteria for manual correction of trait-condition pairs
@@
Appendix D Traits and Conditions
@@
1. Cancer and Oncology-Related Disorders
Liver Cancer
@@ 1. Cancer and Oncology-Related Disorders
Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Susceptibility to Infections, Obstructive sleep apnea, Gender, Anxiety disorder
Kidney Papillary Cell Carcinoma
@@
Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Gender, Susceptibility to Infections, Crohn’s Disease, COVID-19, Obstructive sleep apnea, Anxiety disorder
Kidney Chromophobe
Gender, Age, Hypertension, Obesity, Anxiety disorder
Stomach Cancer
@@
Endometriosis, Gender, Age, Hypertension, Obesity, Anxiety disorder
Bile Duct Cancer
@@
Gender, Age, Obesity, Anxiety disorder, Hypertension
Bladder Cancer
@@
Endometriosis, Glucocorticoid Sensitivity, Gender, Age, Hypertension, Vitamin D Levels, Susceptibility to Infections, Crohn’s Disease, Osteoporosis, Obstructive sleep apnea, COVID-19, Obesity, Alopecia, Anxiety disorder
2. Cardiovascular Diseases
Hypertension
@@ 2. Cardiovascular Diseases
Age, Breast Cancer, Lung Cancer, Gender, Prostate Cancer, Obesity, Endometriosis, Obstructive sleep apnea, Pancreatic Cancer, Vitamin D Levels, Glucocorticoid Sensitivity, COVID-19, Susceptibility to Infections, Bladder Cancer, Kidney Papillary Cell Carcinoma, Osteoporosis, Liver Cancer, Head and Neck Cancer, Esophageal Cancer, Crohn’s Disease, Thyroid Cancer, Colon and Rectal Cancer, Epilepsy, Sjögren’s Syndrome, Anxiety disorder
3. Neurological and Psychiatric Disorders
Multiple Chemical Sensitivity
@@ 3. Neurological and Psychiatric Disorders
Anxiety disorder, Obesity, Age, Hypertension
Anxiety disorder
@@
Obstructive sleep apnea, Gender, Hypertension, Obesity, Age
Amyotrophic Lateral Sclerosis
@@
Age, Hypertension, Obesity, Gender, Anxiety disorder
4. Metabolic and Endocrine Disorders
Glucocorticoid Sensitivity
@@ 4. Metabolic and Endocrine Disorders
Bladder Cancer, Pancreatic Cancer, Endometriosis, Hypertension, Lung Cancer, Breast Cancer, Prostate Cancer, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Obesity, Vitamin D Levels, Crohn’s Disease, Liver Cancer, Osteoporosis, Esophageal Cancer, COVID-19, Obstructive sleep apnea, Susceptibility to Infections, Colon and Rectal Cancer, Anxiety disorder
Osteoporosis
@@
Bone Density, Lung Cancer, Vitamin D Levels, Hypertension, Endometriosis, Breast Cancer, Prostate Cancer, Bladder Cancer, Age, Pancreatic Cancer, Glucocorticoid Sensitivity, Gender, Obstructive sleep apnea, Obesity, Thyroid Cancer, Psoriatic Arthritis, Head and Neck Cancer, Esophageal Cancer, Colon and Rectal Cancer, Crohn’s Disease, Anxiety disorder
Polycystic Kidney Disease
@@
Gender, Hypertension, Obesity, Anxiety disorder
5. Genetic and Developmental Disorders
Multiple Endocrine Neoplasia Type 2
Hypertension, Anxiety disorder, Obesity
Alopecia
@@ 5. Genetic and Developmental Disorders
Psoriatic Arthritis, Endometriosis, Susceptibility to Infections, Crohn’s Disease, Bladder Cancer, Obesity, Hypertension, Anxiety disorder
Intellectual Disability
Age, Obesity, Gender, Hypertension, Anxiety disorder
Craniosynostosis
Obesity, Gender, Age
Brugada Syndrome
Anxiety disorder, Hypertension, Age, Gender, Obesity
Autoinflammatory Disorders
@@
Psoriatic Arthritis, Endometriosis, Hypertension, Obesity, Anxiety disorder
6. Gastrointestinal and Hepatic Disorders
Crohn’s Disease
@@ 6. Gastrointestinal and Hepatic Disorders
Susceptibility to Infections, Pancreatic Cancer, Breast Cancer, Lung Cancer, COVID-19, Bladder Cancer, Age, Glucocorticoid Sensitivity, Prostate Cancer, Endometriosis, Psoriatic Arthritis, Obesity, Celiac Disease, Sjögren’s Syndrome, Vitamin D Levels, Hypertension, Alopecia, Gender, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Head and Neck Cancer, Osteoporosis, Anxiety disorder
Celiac Disease
@@
Crohn’s Disease, Susceptibility to Infections, Sjögren’s Syndrome, Psoriatic Arthritis, COVID-19, Endometriosis, Gender, Obesity, Hypertension, Age, Anxiety disorder
7. Respiratory and Pulmonary Disorders
Obstructive sleep apnea
@@ 7. Respiratory and Pulmonary Disorders
Hypertension, COVID-19, Vitamin D Levels, Obesity, Endometriosis, Prostate Cancer, Bladder Cancer, Osteoporosis, Age, Breast Cancer, Lung Cancer, LDL Cholesterol Levels, Pancreatic Cancer, Susceptibility to Infections, Sjögren’s Syndrome, Thyroid Cancer, Glucocorticoid Sensitivity, Crohn’s Disease, Anxiety disorder, Liver Cancer, Kidney Papillary Cell Carcinoma
COVID-19
@@
Susceptibility to Infections, Obstructive sleep apnea, Hypertension, Sjögren’s Syndrome, Endometriosis, Age, Pancreatic Cancer, Crohn’s Disease, Lung Cancer, Psoriatic Arthritis, Breast Cancer, Gender, Obesity, Prostate Cancer, Bladder Cancer, Vitamin D Levels, Thyroid Cancer, Head and Neck Cancer, Glucocorticoid Sensitivity, Celiac Disease, Kidney Papillary Cell Carcinoma, Anxiety disorder
8. Rheumatological and Musculoskeletal Disorders
Psoriatic Arthritis
@@ 8. Rheumatological and Musculoskeletal Disorders
COVID-19, Sjögren’s Syndrome, Crohn’s Disease, Alopecia, Autoinflammatory Disorders, Susceptibility to Infections, Celiac Disease, Osteoporosis, Obstructive sleep apnea, Gender, Age, Hypertension, Obesity, Anxiety disorder
Sjögren’s Syndrome
@@
Susceptibility to Infections, COVID-19, Psoriatic Arthritis, Endometriosis, Crohn’s Disease, Celiac Disease, Obstructive sleep apnea, Pancreatic Cancer, Hypertension, Thyroid Cancer, Vitamin D Levels, Breast Cancer, Age, Obesity, Anxiety disorder
9. Miscellaneous Traits and Conditions
Endometriosis
@@ 9. Miscellaneous Traits and Conditions
Pancreatic Cancer, Breast Cancer, Lung Cancer, Bladder Cancer, Hypertension, Kidney Papillary Cell Carcinoma, Susceptibility to Infections, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Head and Neck Cancer, Thyroid Cancer, Liver Cancer, COVID-19, Colon and Rectal Cancer, Esophageal Cancer, Obstructive sleep apnea, Osteoporosis, Endometrioid Cancer, Crohn’s Disease, Sjögren’s Syndrome, Alopecia, Stomach Cancer, Autoinflammatory Disorders, Celiac Disease, Anxiety disorder
Susceptibility to Infections
@@
COVID-19, Crohn’s Disease, Endometriosis, Sjögren’s Syndrome, Age, Pancreatic Cancer, Lung Cancer, Hypertension, Breast Cancer, Bladder Cancer, Gender, Prostate Cancer, Obesity, Vitamin D Levels, Thyroid Cancer, Celiac Disease, Psoriatic Arthritis, Alopecia, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Liver Cancer, Glucocorticoid Sensitivity, Esophageal Cancer, Anxiety disorder
Kidney stones
Gender, Hypertension, Obesity, Anxiety disorder
Underweight
Obesity, Gender, Age, Anxiety disorder, Hypertension
Appendix E Details about the data sources

### v3

@@ 5.2.3 Performance of Data Preprocessing
To assess the Data Engineer, we execute gold-standard regression code on TAIS-preprocessed outputs. Table 3 shows that review budget strongly impacts quality: overall F1 improves from 14.41% (0 reviews) to 33.01% (2 reviews), with corresponding SR increasing from 34.83% to 74.02%. Single-step tasks benefit the most (F1 from 21.93% to 47.17%), whereas two-step tasks remain challenging (8.07% to 22.33%), underscoring the difficulty of extracting and harmonizing condition signals from heterogeneous metadata.
Comparing Tables 2 and 3, the dominant bottleneck is preprocessing: even with gold-standard regression, performance remains far below the regression-only setting. Nevertheless, program-and-review substantially reduces errors and almost doubles F1 on two-step tasks, indicating that modest iteration and targeted feedback are highly valuable for data preparation in this domain.
@@ Appendix A Details of the roles
Statistician Statistician agents are assigned the task of performing regression analysis on the preprocessed datasets delivered by the Data Engineer. Its objective is to identify the genes that are predictive of the disease status considering different conditions. To this end, two Statistician agents are designed to perform two categories of regression, i.e. single-step and two-step regression. Both types of Statistician agents will follow the instruction provided by human experts to perform the regression, then analyze the output and interpret and report the results.
@@
Appendix B Criteria for manual correction of trait-condition pairs
@@
Appendix C Traits and Conditions
@@
1. Cancer and Oncology-Related Disorders Liver Cancer Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Susceptibility to Infections, Obstructive sleep apnea, Gender, Anxiety disorder
@@
Kidney Papillary Cell Carcinoma Endometriosis, Age, Hypertension, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Gender, Susceptibility to Infections, Crohn’s Disease, COVID-19, Obstructive sleep apnea, Anxiety disorder
Kidney Chromophobe Gender, Age, Hypertension, Obesity, Anxiety disorder
@@
Stomach Cancer Endometriosis, Gender, Age, Hypertension, Obesity, Anxiety disorder
@@
Bile Duct Cancer Gender, Age, Obesity, Anxiety disorder, Hypertension
@@
Bladder Cancer Endometriosis, Glucocorticoid Sensitivity, Gender, Age, Hypertension, Vitamin D Levels, Susceptibility to Infections, Crohn’s Disease, Osteoporosis, Obstructive sleep apnea, COVID-19, Obesity, Alopecia, Anxiety disorder
@@
2. Cardiovascular Diseases Hypertension Age, Breast Cancer, Lung Cancer, Gender, Prostate Cancer, Obesity, Endometriosis, Obstructive sleep apnea, Pancreatic Cancer, Vitamin D Levels, Glucocorticoid Sensitivity, COVID-19, Susceptibility to Infections, Bladder Cancer, Kidney Papillary Cell Carcinoma, Osteoporosis, Liver Cancer, Head and Neck Cancer, Esophageal Cancer, Crohn’s Disease, Thyroid Cancer, Colon and Rectal Cancer, Epilepsy, Sjögren’s Syndrome, Anxiety disorder
@@
3. Neurological and Psychiatric Disorders Multiple Chemical Sensitivity Anxiety disorder, Obesity, Age, Hypertension
@@
Anxiety disorder Obstructive sleep apnea, Gender, Hypertension, Obesity, Age
@@
Amyotrophic Lateral Sclerosis Age, Hypertension, Obesity, Gender, Anxiety disorder
@@
4. Metabolic and Endocrine Disorders Glucocorticoid Sensitivity Bladder Cancer, Pancreatic Cancer, Endometriosis, Hypertension, Lung Cancer, Breast Cancer, Prostate Cancer, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Obesity, Vitamin D Levels, Crohn’s Disease, Liver Cancer, Osteoporosis, Esophageal Cancer, COVID-19, Obstructive sleep apnea, Susceptibility to Infections, Colon and Rectal Cancer, Anxiety disorder
@@
Osteoporosis Bone Density, Lung Cancer, Vitamin D Levels, Hypertension, Endometriosis, Breast Cancer, Prostate Cancer, Bladder Cancer, Age, Pancreatic Cancer, Glucocorticoid Sensitivity, Gender, Obstructive sleep apnea, Obesity, Thyroid Cancer, Psoriatic Arthritis, Head and Neck Cancer, Esophageal Cancer, Colon and Rectal Cancer, Crohn’s Disease, Anxiety disorder
@@
Polycystic Kidney Disease Gender, Hypertension, Obesity, Anxiety disorder
5. Genetic and Developmental Disorders Multiple Endocrine Neoplasia Type 2 Hypertension, Anxiety disorder, Obesity
@@
Alopecia Psoriatic Arthritis, Endometriosis, Susceptibility to Infections, Crohn’s Disease, Bladder Cancer, Obesity, Hypertension, Anxiety disorder
Intellectual Disability Age, Obesity, Gender, Hypertension, Anxiety disorder
Craniosynostosis Obesity, Gender, Age
Brugada Syndrome Anxiety disorder, Hypertension, Age, Gender, Obesity
@@
Autoinflammatory Disorders Psoriatic Arthritis, Endometriosis, Hypertension, Obesity, Anxiety disorder
@@
6. Gastrointestinal and Hepatic Disorders Crohn’s Disease Susceptibility to Infections, Pancreatic Cancer, Breast Cancer, Lung Cancer, COVID-19, Bladder Cancer, Age, Glucocorticoid Sensitivity, Prostate Cancer, Endometriosis, Psoriatic Arthritis, Obesity, Celiac Disease, Sjögren’s Syndrome, Vitamin D Levels, Hypertension, Alopecia, Gender, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Thyroid Cancer, Head and Neck Cancer, Osteoporosis, Anxiety disorder
@@
Celiac Disease Crohn’s Disease, Susceptibility to Infections, Sjögren’s Syndrome, Psoriatic Arthritis, COVID-19, Endometriosis, Gender, Obesity, Hypertension, Age, Anxiety disorder
@@
7. Respiratory and Pulmonary Disorders Obstructive sleep apnea Hypertension, COVID-19, Vitamin D Levels, Obesity, Endometriosis, Prostate Cancer, Bladder Cancer, Osteoporosis, Age, Breast Cancer, Lung Cancer, LDL Cholesterol Levels, Pancreatic Cancer, Susceptibility to Infections, Sjögren’s Syndrome, Thyroid Cancer, Glucocorticoid Sensitivity, Crohn’s Disease, Anxiety disorder, Liver Cancer, Kidney Papillary Cell Carcinoma
@@
COVID-19 Susceptibility to Infections, Obstructive sleep apnea, Hypertension, Sjögren’s Syndrome, Endometriosis, Age, Pancreatic Cancer, Crohn’s Disease, Lung Cancer, Psoriatic Arthritis, Breast Cancer, Gender, Obesity, Prostate Cancer, Bladder Cancer, Vitamin D Levels, Thyroid Cancer, Head and Neck Cancer, Glucocorticoid Sensitivity, Celiac Disease, Kidney Papillary Cell Carcinoma, Anxiety disorder
@@
8. Rheumatological and Musculoskeletal Disorders Psoriatic Arthritis COVID-19, Sjögren’s Syndrome, Crohn’s Disease, Alopecia, Autoinflammatory Disorders, Susceptibility to Infections, Celiac Disease, Osteoporosis, Obstructive sleep apnea, Gender, Age, Hypertension, Obesity, Anxiety disorder
@@
Sjögren’s Syndrome Susceptibility to Infections, COVID-19, Psoriatic Arthritis, Endometriosis, Crohn’s Disease, Celiac Disease, Obstructive sleep apnea, Pancreatic Cancer, Hypertension, Thyroid Cancer, Vitamin D Levels, Breast Cancer, Age, Obesity, Anxiety disorder
@@
9. Miscellaneous Traits and Conditions Endometriosis Pancreatic Cancer, Breast Cancer, Lung Cancer, Bladder Cancer, Hypertension, Kidney Papillary Cell Carcinoma, Susceptibility to Infections, Glucocorticoid Sensitivity, Vitamin D Levels, Obesity, Head and Neck Cancer, Thyroid Cancer, Liver Cancer, COVID-19, Colon and Rectal Cancer, Esophageal Cancer, Obstructive sleep apnea, Osteoporosis, Endometrioid Cancer, Crohn’s Disease, Sjögren’s Syndrome, Alopecia, Stomach Cancer, Autoinflammatory Disorders, Celiac Disease, Anxiety disorder
@@
Susceptibility to Infections COVID-19, Crohn’s Disease, Endometriosis, Sjögren’s Syndrome, Age, Pancreatic Cancer, Lung Cancer, Hypertension, Breast Cancer, Bladder Cancer, Gender, Prostate Cancer, Obesity, Vitamin D Levels, Thyroid Cancer, Celiac Disease, Psoriatic Arthritis, Alopecia, Obstructive sleep apnea, Kidney Papillary Cell Carcinoma, Liver Cancer, Glucocorticoid Sensitivity, Esophageal Cancer, Anxiety disorder
Kidney stones Gender, Hypertension, Obesity, Anxiety disorder
Underweight Obesity, Gender, Age, Anxiety disorder, Hypertension
Appendix D Details about the data sources